# KG-QAGen: A Knowledge-Graph-Based Framework for Systematic Question Generation and Long-Context LLM Evaluation

### v1

@@ D.1 Prompting Strategy
Batch QA Prompt [SYSTEM INPUT]
You are a financial expert, and your task is to answer the questions given to you in batches about the provided credit agreement. If you believe the answer is not present in the agreement, say ’Not found’.
[EXPECTED OUTPUT]
Respond ONLY with valid JSON, nothing else. See the example below.
The given document:
Tesla, Inc. is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning. Elon Musk became the largest investor and later CEO.
The given questions:
Q1: Who founded Tesla?
Q2: What year was Tesla founded?
The expected output:
    "answers": [               {"question_index": 1,               "answer": "Martin Eberhard,                                 Marc Tarpenning"},               {"question_index": 2,               "answer": "2003"}        ] }
[USER INPUT]
{Document Context}
[QUESTIONS]
Q1: {question}
Q2: {question}
@@
Merging Prompt [SYSTEM INPUT]
You are a financial expert, and your task is to combine or merge the provided partial answers, coming from different chunks of a credit agreement, into a single final answer for each of the questions given to you. If you believe the answer is not present in the agreement, say ’Not found’.
[EXPECTED OUTPUT]
The given document:
Tesla, Inc. is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning. Elon Musk became the largest investor and later CEO.
The given questions:
Q1: Who founded Tesla?
Q2: What year was Tesla founded?
The expected output:
    "answers": [               {"question_index": 1,               "answer": "Martin Eberhard,                                 Marc Tarpenning"},               {"question_index": 2,               "answer": "2003"}        ] }
[USER INPUT]
Chunk 1 partial answer JSON: {ans 1}
Chunk 2 partial answer JSON: {ans 2}
[QUESTIONS]
Q1: {question}
Q2: {question}
@@
We implement a two-stage prompting pipeline for QA inference over long documents. In the first stage, the document is either processed in full or split into chunks (depending on the context limit of the LLM), and each chunk is paired with a batch of up to 50 questions. We then apply a one-shot, instruction-based prompt (Figure 7(a)) that instructs the LLM to act as a financial expert and return an array of answers – either grounded in the document or explicitly stating “Not found" – in valid JSON format.
The expected format is enforced both through the prompt instructions and through structured parsing logic. The parser first attempts to extract a JSON object from the model’s response, supporting both raw JSON and code-fenced output. If parsing fails, the system falls back to extracting answers from list-formatted outputs or reattempts the query, up to a maximum number of retries. All models are queried with temperature=0.0 for determinism.
Regarding resources, as mentioned in 4.1, for open-source models we are using Together.AI for some of the models. For the other open-source models, we used 1 H200 GPU.
D.2 Evaluation Metrics
We employ four complementary metrics to evaluate model responses.
  • Word-Level F1 Score  Token-level precision and recall against the gold answer, ensures strict factual correctness in extractive question answering.
  • Cosine Similarity  Evaluates the semantic alignment, providing robustness to minor changes.
@@ D.2 Evaluation Metrics
D.3 Comparison with Oracle and RAG
We compare three retrieval settings for answering questions from credit agreements: (1) Oracle, which uses the pieces of text containing the answer; (2) RAG (Retrieval-Augmented Generation), which selects top passages using a retriever; and (3) Full, where the entire agreement is provided to the model. In table 13, we observe that RAG performs drastically worse than the other two settings across all difficulty levels, which highlights the incompatibility of standard retrieval techniques with dense, cross-referenced financial documents, often requiring the aggregation of logically related but spatially distant elements.
Interestingly, the comparison between Oracle and Full Document settings reveals a non-trivial pattern. For easy look-up questions, the Oracle setting performs better, which aligns with expectations since such questions benefit from isolating the minimal, most relevant snippet. However, for medium and hard questions, the Full Document consistently outperforms the Oracle setting. This suggests that isolating the “correct" snippet removes necessary contextual cues that models need for logical, multi-hop inference. These findings reinforce the value of credit agreements as a testbed for long-context LLMs, and demonstrate how our framework’s question complexity help uncover the precise scenarios where models struggle, even when given the exact ground-truth passage (see Section 5 and E for the detailed analysis).
Table 13: Performance comparison across three retrieval settings – Oracle (gold snippet only), RAG (retriever-selected passages), and Full (entire credit agreement) – for each question complexity group.
Easy Medium Hard
Full 0.8542 0.1252 0.8507 4.4893 0.6936 0.3417 0.6932 3.8983 0.1172 0.7238 0.0608 2.2838
Oracle 0.9113 0.0673 0.9076 4.6950 0.6335 0.4075 0.6360 4.0508 0.1128 0.8067 0.0736 1.9189
RAG 0.1481 0.7341 0.1473 1.6099 0.1277 0.7543 0.1251 1.5254 0.0191 0.8450 0.0107 1.2027
D.4 Human Evaluation Metrics
  • 1   response is irrelevant or does not match the actual answer;
  • 2   response is somewhat relevant but does not match the actual answer;
  • 3   response shows slight similarity but lacks accuracy;
  • 4   response is largely relevant and accurate;
  • 5   response is a perfect match to the actual answer.
Table 14: Kendall’s Tau and Pearson Correlation Tables.
@@ D.4 Human Evaluation Metrics
Pearson Correlation Matrix
F1 Score 1.000 0.933 -0.894 0.922 0.941
Cosine Similarity 0.933 1.000 -0.910 0.935 0.922
Edit Distance Sim -0.894 -0.910 1.000 -0.844 -0.862
LLM-as-a-Judge 0.922 0.935 -0.844 1.000 0.947
Avg LLM-Human Rating 0.941 0.922 -0.862 0.947 1.000
Figure 8: Kendall’s Tau and Pearson Correlation Matrices.
@@ Appendix E Error Analysis
Below are detailed examples of errors from each of the aforementioned error categories.
@@
Error Explanation Misinterpretation of Semantics. The LLM misinterprets the semantic intent of the question. While City is factually accurate, the question requires identifying the Location Type of "Dubai, UAE" in the context of the credit agreement. The expected answer, zonal office, reflects a more precise understanding of this context.
@@
Error Explanation Implicit Information Gaps. The LLM fails to infer implicit information due to structural complexity in the document. In cases involving multiple signatories or nested signature panels, human readers can intuitively associate individuals with their respective organizations based on layout and context. However, this relationship is not explicitly stated, and the LLM, lacking this inferential reasoning, incorrectly attributes Hamish Sandhu to the wrong organization.
@@
Error Explanation Set Operation Failures. The LLM struggles with this query due to the complexity introduced by multiple set operations, in this case, requiring a comparison of roles across three individuals. Alexander Wu is the SVP and General Manager, while Chwan-Ming Ho and Edward Chen are both VPs and General Managers (see figure  9). While "SVP" is correct, the inclusion of "General Manager" indicates a failure to correctly subtract shared titles.
@@
Error Explanation Long-Context Retrieval Limitations. The model fails to retrieve the correct role because the organizational label "subsidiary" is not explicitly stated in a natural sentence. Instead, Celanese Chemicals, Inc. appears beneath a heading or clause indicating subsidiaries, often following the signature panel. This implicit semantic link requires structural and hierarchical reasoning that current LLMs struggle with—particularly when role information is implied by layout rather than expressed textually.
@@
Appendix F Dataset Release and Leaderboard Setup
Table 16: Dataset statistics of the constructed KG-QAGen-D dataset.

### v2

@@ D.1 Prompting Strategy
We implement a two-stage prompting pipeline for QA inference over long documents. In the first stage, the document is either processed in full or split into chunks (depending on the context limit of the LLM), and each chunk is paired with a batch of up to 50 questions. We then apply a one-shot, instruction-based prompt (Figure 7(a)) that instructs the LLM to act as a financial expert and return an array of answers – either grounded in the document or explicitly stating “Not found" – in valid JSON format. The effect of chunk-based evaluation is discussed in Appendix D.5.
The expected format is enforced both through the prompt instructions and through structured parsing logic. The parser first attempts to extract a JSON object from the model’s response, supporting both raw JSON and code-fenced output. If parsing fails, the system falls back to extracting answers from list-formatted outputs or reattempts the query, up to a maximum number of retries. All models are queried with temperature=0.0.
Regarding resources, as mentioned in Section 3.1, for open-weight models we are using Together.AI for some of the models. For the other open-weight models, we used 1 H200 GPU.
D.2 Evaluation Metrics
We employ four complementary metrics to evaluate model responses (Table 12).
  • Word-Level F1 Score  Token-level precision and recall against the gold answer, ensures strict factual correctness in question answering.
  • Cosine Similarity  Evaluates the semantic alignment, with robustness to minor changes.
D.3 Human Evaluation Metrics
  • 1   response is irrelevant or does not match the actual answer;
  • 2   response is somewhat relevant but does not match the actual answer;
  • 3   response shows slight similarity but lacks accuracy;
  • 4   response is largely relevant and accurate;
  • 5   response is a perfect match to the actual answer.
Metric prioritization and interpretation
Based on the observed correlations with human judgments, we recommend interpreting LLM-as-a-Judge as the primary indicator of semantic correctness, particularly for higher-complexity questions where answers may be paraphrased, partially implicit, or expressed with variable surface forms. F1 score should be viewed as a complementary metric that captures extractive fidelity and coverage, and is most informative when exact grounding in the document is required. Discrepancies between the two metrics are therefore not treated as inconsistencies but as signals of different model behaviors (e.g., cautious but semantically accurate answers versus broader but partially grounded ones), as discussed in Section 3.3. For practical use, we recommend reporting both metrics jointly: LLM-as-a-Judge to assess answer correctness from a human-aligned perspective, and F1 to diagnose retrieval completeness and over-/under-generation. The remaining metrics (Edit Distance and Cosine Similarity) provide additional diagnostic insight into surface-level mismatch and semantic drift but are not intended to be used in isolation for model ranking.
Batch QA Prompt [SYSTEM INPUT] You are a financial expert, and your task is to answer the questions given to you in batches about the provided credit agreement. If you believe the answer is not present in the agreement, say ’Not found’. [EXPECTED OUTPUT] Respond ONLY with valid JSON, nothing else. See the example below. The given document: Tesla, Inc. is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning. Elon Musk became the largest investor and later CEO. The given questions: Q1: Who founded Tesla? Q2: What year was Tesla founded? The expected output: { "answers": [       {"question_index": 1,       "answer": "Martin Eberhard,               Marc Tarpenning"},       {"question_index": 2,       "answer": "2003"}    ] } [USER INPUT] {Document Context} [QUESTIONS] Q1: {question} Q2: {question}
@@ D.3 Human Evaluation Metrics
Merging Prompt [SYSTEM INPUT] You are a financial expert, and your task is to combine or merge the provided partial answers, coming from different chunks of a credit agreement, into a single final answer for each of the questions given to you. If you believe the answer is not present in the agreement, say ’Not found’. [EXPECTED OUTPUT] The given document: Tesla, Inc. is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning. Elon Musk became the largest investor and later CEO. The given questions: Q1: Who founded Tesla? Q2: What year was Tesla founded? The expected output: { "answers": [       {"question_index": 1,       "answer": "Martin Eberhard,               Marc Tarpenning"},       {"question_index": 2,       "answer": "2003"}    ] } [USER INPUT] Chunk 1 partial answer JSON: {ans 1} Chunk 2 partial answer JSON: {ans 2} [QUESTIONS] Q1: {question} Q2: {question}
@@
Figure 8: Pearson and Kendall’s Tau Correlation Matrices.
Table 13: Pearson and Kendall’s Tau Correlation Tables.
Pearson Correlation Matrix
F1 Score 1.000 0.933 -0.894 0.922 0.941
Cosine Similarity 0.933 1.000 -0.910 0.935 0.922
Edit Distance Sim -0.894 -0.910 1.000 -0.844 -0.862
LLM-as-a-Judge 0.922 0.935 -0.844 1.000 0.947
Avg LLM-Human Rating 0.941 0.922 -0.862 0.947 1.000
@@
D.4 Cost-Constrained Models
For the cost-constrained models, we used stratified sampling over our question templates. Because each template corresponds to a unique point in the (H, P, #SO) space, sampling proportionally over templates preserves the difficulty distribution, the multi-hop vs single-hop ratio, and the relative frequency of set-operation queries. Using DeepSeek-V3, the strongest open-weight model evaluated on the full dataset, we have verified that performance on the sampled subset correlates strongly with full-set performance: 0.87 for F-score, 0.84 for edit distance, and 0.88 for cosine similarity.
D.5 Chunk-Based Evaluation
As seen in Table 14, the effect of chunking becomes noticeable as the question complexity grows. The performance increase is driven mainly by a reduction in abstentions, with models more often returning explicit answers instead of ‘‘Not Found’’. However, even for medium questions nearly half is still unanswerable without chunking: it does not reflect improved reasoning or extraction quality, which means chunking does not affect our error analysis or the conclusions drawn in the paper.
Table 14: Evaluation of Gemini-2.0-Flash with chunking (as in the main paper) and without chunking. The context length of this model is large enough to incorporate an entire credit agreement in one prompt.
Strategy F1 $\uparrow$ Edit $\downarrow$ Cos $\uparrow$ LLM-J $\uparrow$
Easy
With 0.6748 0.2679 0.6650 4.0000
Without 0.7105 0.2987 0.6949 3.9269
Medium
With 0.3294 0.6280 0.3121 2.4820
Without 0.5558 0.4051 0.5452 3.4344
Hard
With 0.1172 0.7238 0.0608 2.2838
Without 0.2565 0.6310 0.2504 2.1088
D.6 Single Question Evaluation
We note that the absolute metric values in Table 15 are not representative of task complexity, as a sampling of 20 questions per complexity level was required due to the high cost of single-question evaluation for long-context documents. Examining relative differences, the single-question evaluation provides marginal gains for easy and medium questions, but not for hard ones. These variations are minor and leave the overall performance patterns across complexity levels unchanged.
Table 15: Evaluation of Gemini-2.0-Flash in batches of 50 questions (as in the main paper) and with a simple questions.
Difficulty F1 $\uparrow$ Edit $\downarrow$ Cos $\uparrow$ LLM-J $\uparrow$
Easy
Batch 0.5277 0.3773 0.5249 3.7500
Single 0.5719 0.3573 0.5688 3.8000
Medium
Batch 0.3980 0.4220 0.3958 3.2040
Single 0.4451 0.3695 0.4427 3.5510
Hard
Batch 0.5333 0.4223 0.5333 3.1333
Single 0.4971 0.4367 0.4938 3.0666
D.7 Human Performance Evaluation
To establish a reference point for model comparison, we conducted a human evaluation on a balanced subset of 100 questions (20 per difficulty level). Three independent human evaluators participated in this study: one annotator who was fully familiar with the documents, and two participants with general financial knowledge but no prior exposure to the documents. Their responses were normalized and compared against ground truth using three similarity metrics: word level F1, edit distance and cosine similarity. Unlike model performance, which typically declines as task complexity increases, human performance remained relatively stable across levels, as seen in Table 16. This suggests that humans effectively leverage contextual reasoning and background knowledge to maintain accuracy even in multi-hop or compositional settings, capabilities that current models still struggle with.
Table 16: Mean human performance across question levels.
Level F1 Edit Dist. Cosine Sim.
1 0.846 0.142 0.834
2 0.778 0.264 0.750
3 0.780 0.235 0.756
4 0.803 0.244 0.795
5 0.935 0.101 0.928
Table 17: Performance comparison across four retrieval settings – Full (entire credit agreement), Oracle (gold snippet only), RAG and dynamic RAG (retriever-selected passages) – for each question complexity group.
Easy Medium Hard
Full 0.6747 0.2678 0.6650 4.000 0.3294 0.6279 0.3121 2.4820 0.1172 0.7237 0.0607 2.2837
Oracle 0.7134 0.2588 0.7021 3.9862 0.3294 0.5986 0.3201 2.6497 0.1127 0.8066 0.0735 1.9189
RAG 0.1579 0.7364 0.1546 1.6931 0.0695 0.8070 0.0651 1.4131 0.0190 0.8449 0.0107 1.2027
D-RAG 0.1671 0.7304 0.1631 1.7185 0.1118 0.7693 0.1087 1.5566 0.0107 0.8431 0.0069 1.1088
D.8 Comparison with Oracle and RAG
We compare four retrieval settings for answering questions from credit agreements: (1) Full, where the entire agreement is provided to the model, (2) Oracle, which uses the pieces of text containing the answer; (3) RAG (Retrieval-Augmented Generation), which selects top passages using a retriever; and (4) Dynamic RAG, which iteratively decomposes the question into sub-queries, retrieves relevant passages across multiple steps, and aggregates the retrieved evidence before generating the final answer. In Table 17, we observe that RAG, both simple and dynamic, performs drastically worse than the other two settings across all difficulty levels, which highlights the incompatibility of standard retrieval techniques with dense, cross-referenced financial documents, often requiring the aggregation of logically related but spatially distant elements.
Interestingly, the comparison between Oracle and Full Document settings reveals a non-trivial pattern. For easy look-up and for medium questions, the Oracle setting performs better, as expected, since such questions benefit from isolating the minimal, most relevant snippet. However, for hard questions, the Full Document consistently outperforms the Oracle setting. This suggests that isolating the “correct” snippet removes necessary contextual cues that models need for reasoning, which is detailed in the Implicit Information Gaps paragraph of our Error Analysis (Section 4.2). These findings reinforce the value of credit agreements as a testbed for long-context LLMs, and demonstrate how our framework’s question complexity help uncover the precise scenarios where models struggle, even when given the exact ground-truth passage (see Section 4 and E for the detailed analysis).
@@ Appendix E Error Analysis
Detailed examples of errors from each of the aforementioned categories are shown in Table 18 and Figure 9. Aggregate trends across error types and complexity are shown in Figures 10 and 11.
@@
Error Explanation Misinterpretation of Semantics. The LLM misinterprets the semantic intent of the question. While City is factually accurate, the question requires identifying the Location Type of "Dubai, UAE" in the context of the credit agreement. The expected answer, zonal office, reflects a more precise understanding.
@@
Error Explanation Implicit Information Gaps. The LLM fails to infer implicit information due to structural complexity. Human readers can associate individuals with their organizations based on layout and context. The LLM incorrectly attributes Hamish Sandhu to the wrong organization.
@@
Error Explanation Set Operation Failures. The LLM struggles with multiple set operations. Alexander Wu is SVP and General Manager, while Chwan-Ming Ho and Edward Chen are both VPs and General Managers. The extra "General Manager" indicates a failure in subtracting shared titles.
@@
Error Explanation Long-Context Retrieval Limitations. The model fails because "subsidiary" is not explicitly stated in a natural sentence. Instead, Celanese Chemicals, Inc. appears under headings or clauses indicating subsidiaries, requiring structural and hierarchical reasoning.
@@
Appendix F Ablation Study on Medical Documents
To illustrate the generalizability and applicability of our framework to other domain, we conducted a minor ablation study on the publicly available [^8] https://clinicalinfo.hiv.gov/en/guidelines/hiv-clinical-guidelines-adult-and-adolescent-opportunistic-infections/whats-new "Guidelines for the Prevention and Treatment of Opportunistic Infections in Adults and Adolescents With HIV". We extracted the introductory part, the first 10 descriptions of diseases, and the conclusive part, which results in a 225-page document. We annotated the document according to the schema in Figure 12, which allowed us to extract questions of various complexity levels over the same 3 dimensions as in the main study (see Tables 20-22 for the full list of question templates).
Figure 12: Annotation schema of our ablation study.
We ran our experiments with the best-performing open-weight model, DeepSeek-V3. In Table 19, we observe a clear performance decrease as the question complexity increases, which illustrates that our framework is generalizable and applicable to other domains.
Table 19: Evaluation of DeepSeek-V3 on the ablation study dataset.
Difficulty F1 $\uparrow$ Edit $\downarrow$ Cos $\uparrow$ LLM-J $\uparrow$
Easy 0.2585 0.6169 0.2522 2.2249
Medium 0.0753 0.7512 0.0662 1.5587
Hard 0.0 0.8417 0.0 1.0681
Table 20: Ablation 1-hop Question Templates
Template P H #SO
What pathogen is transmitted by [Vector]? [if one] 0 1 0
What vector transmits [Pathogen]? [if one] 0 1 0
What pathogen causes [Disease]? [if one] 0 1 0
What disease is caused by [Pathogen]? [if one] 0 1 0
What disease is treated by [Medication]? [if one] 0 1 0
What medication treats [Disease]? [if one] 0 1 0
What pathogens are transmitted by [Vector]? 1 1 0
What vectors transmit [Pathogen]? 1 1 0
What pathogens cause [Disease]? 1 1 0
What diseases are caused by [Pathogen]? 1 1 0
What diseases are treated by [Medication]? 1 1 0
What medications treat [Disease]? 1 1 0
What pathogen is transmitted by [Vector 1] and [Vector 2]? [if one] 0 1 1
What vector transmits [Pathogen 1] and [Pathogen 2]? [if one] 0 1 1
What pathogen causes [Disease 1] and [Disease 2]? [if one] 0 1 1
What disease is caused by [Pathogen 1] and [Pathogen 2]? [if one] 0 1 1
What disease is treated by [Medication 1] and [Medication 2]? [if one] 0 1 1
What medication treats [Disease 1] and [Disease 2]? [if one] 0 1 1
What pathogens are transmitted by [Vector 1] and [Vector 2]? 1 1 1
What vectors transmit [Pathogen 1] and [Pathogen 2]? 1 1 1
What pathogens cause [Disease 1] and [Disease 2]? 1 1 1
What diseases are caused by [Pathogen 1] and [Pathogen 2]? 1 1 1
What diseases are treated by [Medication 1] and [Medication 2]? 1 1 1
What medications treat [Disease 1] and [Disease 2]? 1 1 1
What pathogen is transmitted by [Vector 1] but not [Vector 2]? [if one] 0 1 2
What vector transmits [Pathogen 1] but not [Pathogen 2]? [if one] 0 1 2
What pathogen causes [Disease 1] but not [Disease 2]? [if one] 0 1 2
What disease is caused by [Pathogen 1] but not [Pathogen 2]? [if one] 0 1 2
What disease is treated by [Medication 1] but not [Medication 2]? [if one] 0 1 2
What medication treats [Disease 1] but not [Disease 2]? [if one] 0 1 2
What pathogens are transmitted by [Vector 1] but not [Vector 2]? 1 1 2
What vectors transmit [Pathogen 1] but not [Pathogen 2]? 1 1 2
What pathogens cause [Disease 1] but not [Disease 2]? 1 1 2
What diseases are caused by [Pathogen 1] but not [Pathogen 2]? 1 1 2
What diseases are treated by [Medication 1] but not [Medication 2]? 1 1 2
What medications treat [Disease 1] but not [Disease 2]? 1 1 2
Table 21: Ablation 2-hop Question Templates
Template P H #SO
What vector transmits a pathogen which causes [Disease]? [if one] 0 2 0
What disease is caused by a pathogen which is transmitted by [Vector]? [if one] 0 2 0
What medication treats a disease which is caused by [Pathogen]? [if one] 0 2 0
What pathogen causes a disease which is treated by [Medication]? [if one] 0 2 0
What vectors transmit a pathogen which causes [Disease]? 1 2 0
What diseases are caused by a pathogen which is transmitted by [Vector]? 1 2 0
What medications treat a disease which is caused by [Pathogen]? 1 2 0
What pathogens cause a disease which is treated by [Medication]? 1 2 0
What vector transmits a pathogen which causes [Disease 1] and [Disease 2]? [if one] 0 2 1
What disease is caused by a pathogen which is transmitted by [Vector 1] and [Vector 2]? [if one] 0 2 1
What medication treats a disease which is caused by [Pathogen 1] and [Pathogen 2]? [if one] 0 2 1
What pathogen causes a disease which is treated by [Medication 1] and [Medication 2]? [if one] 0 2 1
What vectors transmit a pathogen which causes [Disease 1] and [Disease 2]? 1 2 1
What diseases are caused by a pathogen which is transmitted by [Vector 1] and [Vector 2]? 1 2 1
What medications treat a disease which is caused by [Pathogen 1] and [Pathogen 2]? 1 2 1
What pathogens cause a disease which is treated by [Medication 1] and [Medication 2]? 1 2 1
What vector transmits a pathogen which causes [Disease 1] but not [Disease 2]? [if one] 0 2 2
What disease is caused by a pathogen which is transmitted by [Vector 1] but not [Vector 2]? [if one] 0 2 2
What medication treats a disease which is caused by [Pathogen 1] but not [Pathogen 2]? [if one] 0 2 2
What pathogen causes a disease which is treated by [Medication 1] but not [Medication 2]? [if one] 0 2 2
What vectors transmit a pathogen which causes [Disease 1] but not [Disease 2]? 1 2 2
What diseases are caused by a pathogen which is transmitted by [Vector 1] but not [Vector 2]? 1 2 2
What medications treat a disease which is caused by [Pathogen 1] but not [Pathogen 2]? 1 2 2
What pathogens cause a disease which is treated by [Medication 1] but not [Medication 2]? 1 2 2
Table 22: Ablation 3-hop Question Templates
Template P H #SO
What vector transmits a pathogen which causes a disease which is treated by [Medication]? [if one] 0 3 0
What medication treats a disease which is caused by a pathogen which is transmitted by [Vector]? [if one] 0 3 0
What vectors transmit a pathogen which causes a disease which is treated by [Medication]? 1 3 0
What medications treat a disease which is caused by a pathogen which is transmitted by [Vector]? 1 3 0
What vector transmits a pathogen which causes a disease which is treated by [Medication 1] and [Medication 2]? [if one] 0 3 1
What medication treats a disease which is caused by a pathogen which is transmitted by [Vector 1] and [Vector 2]? [if one] 0 3 1
What vectors transmit a pathogen which causes a disease which is treated by [Medication 1] and [Medication 2]? 1 3 1
What medications treat a disease which is caused by a pathogen which is transmitted by [Vector 1] and [Vector 2]? 1 3 1
What vector transmits a pathogen which causes a disease which is treated by [Medication 1] but not [Medication 2]? [if one] 0 3 2
What medication treats a disease which is caused by a pathogen which is transmitted by [Vector 1] but not [Vector 2]? [if one] 0 3 2
What vectors transmit a pathogen which causes a disease which is treated by [Medication 1] but not [Medication 2]? 1 3 2
What medications treat a disease which is caused by a pathogen which is transmitted by [Vector 1] but not [Vector 2]? 1 3 2
Appendix G Dataset Release and Leaderboard Setup
To facilitate reproducibility and future research (Tatarinov et al., 2025), we are publicly releasing (under CC-BY NC ND 4.0 license) QA pairs for 40 documents out of the total 170 as the development set, allowing researchers to validate model performance under controlled conditions. The remaining QA pairs serve as the test set, which is not publicly available to prevent data contamination (Sainz et al., 2023). For this test set, we are making the questions public, without the ground-truth answers. The distribution of the test and dev sets is provided in Table 23. In addition, we are hosting a leaderboard [^9] https://huggingface.co/spaces/gtfintechlab/KG-MuLQA-D-Leaderboard, following standard practices in recent LLM evaluation benchmarks (Galarnyk et al., 2025; Yue et al., 2024; Zhao et al., 2024a; Lu et al., 2023).
Table 23: Dataset statistics of the constructed KG-MuLQA-D dataset.